OpenStudy (anonymous):

If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. What is the numerator of the fraction? 2 9 16

OpenStudy (vickyisthesmartone123):

So we have : nd−7=1 n = d-7 d = n + 7 ----------(1)

OpenStudy (vickyisthesmartone123):

Now... since n+3d=34 4(n+3)=3d 4n+12=3d Put d = n+7 4n + 12 = 3(n+7) 4n+ 12 = 3n +21 n = 9 So we have numerator = 9 and d = n + 7 = 16 hence numerator = 9 and denominator = 16

OpenStudy (vickyisthesmartone123):

Do you get it?